OpenStudy (babyslapmafro):

Please help me find the following limit using L'Hopital's rule...(click to see)

OpenStudy (babyslapmafro):

\[\lim_{x \rightarrow \infty }xsin(\frac{ \pi}{ x } )\]

OpenStudy (babyslapmafro):

Is there a more simple route than converting to x/csc(pi/x)

OpenStudy (anonymous):

maybe convert it to \[\frac{\sin(\frac{\pi}{x})}{\frac{1}{x}}\]

OpenStudy (babyslapmafro):

ok ill try that